A professional photographer who specializes in wedding-related activities paid $43,500 for equipment that will have a $2000 salvage value after 5 years. He estimates that his costs associated with each event amount to $65 per day. If he charges $300 per day for his services, how many days per year must he be employed in order to break even at an interest rate of 10% per year? The number of days required for break even is determined to be per year.
Answer #1

Initial investment = $43500

Salvage value = $2000

Time = 5 years

R = 10%

PW of the investment = -43500 + 2000/1.1^5 = -$42258.2

Let, uniform annual investment = P

Then,

P = -42258.2/((1-1/1.1^5)/.1)

P = -$11147.6

So, uniform annual investment or annual fixed cost is $11147.6.

revenue per event per day = $300

Variable cost per event per day = $65

Hence,

Break-even number of days per year = 11147.6/(300-65)

Break-even number of days per year = 47.44 days